# Bridging Today and the Future of Humanity: AI Safety in 2024 and Beyond

### v3

@@ 7 Conclusion
This paper presents a blueprint for an advanced human society and leverages this forward-looking vision to guide today’s AI safety efforts. Through the blueprint, it becomes clear that Artificial General Intelligence (AGI) is not the ultimate goal of AI development. Instead, the true vision lies in the Internet of Everything, a deeply interconnected world where intelligent systems seamlessly integrate into every aspect of life. While AGI is a popular topic discussed more among people in the computer science field today, a more advanced world demands interdisciplinary collaborations across various fields. Regarding AGI as the ultimate goal might limit our creativity in the 2020s, trapping us in “local optima” and potentially causing us to overlook the real challenges.
What will this future world be like? Will it bring more happiness to human-beings? The answer is uncertain. Advanced technologies may take on more duties of daily life, leaving humans with fewer tasks and more predictability in everyday life. On the other hand, with many everyday problems being solved by advanced technologies, people may have more time for self-reflection and personal growth, which may lead to another form of “happiness”.
As we stand at a pivotal moment in human history, our efforts become not only a technological endeavor but also a profound exploration of human potential and the boundaries of human capability. In any case, the journey toward this advanced future is fascinating, and every human being will look forward to it.

### v4

@@ 7 Conclusion
This paper presents a blueprint for an advanced human society and leverages this forward-looking vision to guide today’s AI safety efforts. Through the blueprint, it becomes clear that Artificial General Intelligence (AGI) is not the ultimate goal of AI development. Instead, the true vision lies in the Internet of Everything, a deeply interconnected world where intelligent systems seamlessly integrate into daily life. While AGI is a popular topic discussed more among people in the computer science field today, a more advanced world demands interdisciplinary collaborations across various fields. Regarding AGI as the ultimate goal might limit our creativity in the 2020s, trapping us in “local optima” and potentially causing us to overlook the real challenges.
What will this future world be like? Will it bring more happiness to human-beings? The answer is uncertain. Advanced technologies may take on more duties of daily life, leaving humans with fewer tasks and more predictability. On the other hand, with more daily problems being solved by technologies, people may have more time for self-reflection and personal growth, which may lead to another form of “happiness”.
As we stand at a pivotal moment in history, our efforts become more than an incremental technological endeavor, but a profound exploration of human potential and the boundaries of human capability. In any case, the journey toward this advanced future is fascinating, and every human being will look forward to it.